# Leveraging Skills from Unlabeled Prior Data for Efficient Online Exploration

### v1

@@ 5.3 Can we leverage unsupervised trajectory skills to accelerate online learning?
Figure 3 shows the aggregated performance of our approach on all three domains. Overall, our method outperforms all prior methods. Both HILP-based methods (Online with HILP Skills and HILP with Offline Data) achieve a good initial return on the Kitchen domain, but improve very slowly during online training. The Online with Trajectory Skills baseline also consistently performs worse than our method across all three domains, which demonstrates the importance of using prior data for online learning of the high-level policy, since that is only difference between this baseline and Ours. ExPLORe uses offline data during online learning, but does not pretrain skills, leading to slower learning on all three environments and difficulty achieving any significant return on Visual AntMaze or Kitchen. We also report the performance on individual AntMaze mazes and Kitchen tasks in Figure 4, and observe that our method outperforms the baselines more on harder environments. For the AntMaze domain, as the maze size increases, the two best baselines, Online w/ Trajectory Skills and ExPLORe, perform worse, while Ours continues to achieve a similar level of performance. For the Kitchen domain, on the more challenging kitchen-mixed and kitchen-partial tasks, the improvement of Ours over Online w/ Trajectory Skills is also larger. These experiments suggest that pretraining skills and the ability to leverage the prior data are both crucial for achieving efficient online learning, especially in more challenging environments. For the visual domain, we additionally perform an ablation study to assess the importance of the ICVF pre-trained representation, which we include in Appendix D. While ICVF combines synergistically with our method to further accelerate learning and exploration, initializing RND image encoder weights using ICVF is not critical to its success. Even though we have demonstrated that our method is able to achieve higher success rate faster than prior works, it is still not clear if our method can actually lead to better exploration (instead of simply learning the high-level policy better). In the following two sections, we study the exploration aspect in isolation in the AntMaze domain.
@@ 5.4 Is our method able to find goals faster than prior methods?
Table 1 shows the average number of online environment interaction steps for the agent to reach the goal for each of the four goals in each of the three maze layouts. Such a metric allows us to assess how efficiently the agent explores in the maze whereas the success rate metric only measures how good the agent is at reaching the desired goal. It is possible for an agent to be good at exploration, but bad at reaching goals consistently and vice-versa. Table 1 shows that our method’s average first goal time outperforms or matches every baseline on each of the 12 goals, which confirms that our method not only learns faster, but does so by exploring more efficiently.
Maze Layout Goal Location Methods without Pretraining Methods with Pretraining Online + RND ExPLORe Diffusion BC w/ JSRL Online w/ Trajectory Skills Online w/ HILP Skills HILP w/ Offline Data Ours Medium Top Left 71 $\pm$ 5.0 27 $\pm$ 3.2 60 $\pm$ 8.1 21 $\pm$ 4.1 120 $\pm$ 47 27 $\pm$ 6.2 14 $\pm$ 3.1 Top Right 100 $\pm$ 16 29 $\pm$ 2.8 85 $\pm$ 19 76 $\pm$ 26 160 $\pm$ 40 72 $\pm$ 36 22 $\pm$ 3.2 Bottom Right 230 $\pm$ 38 35 $\pm$ 4.9 99 $\pm$ 15 77 $\pm$ 34 300 $\pm$ 0 270 $\pm$ 33 22 $\pm$ 4.4 Center 210 $\pm$ 32 71 $\pm$ 8.0 260 $\pm$ 28 26 $\pm$ 3.4 260 $\pm$ 28 300 $\pm$ 0.0 18 $\pm$ 1.7 Aggregated 150 $\pm$ 14 40 $\pm$ 2.0 130 $\pm$ 10 50 $\pm$ 11 210 $\pm$ 17 170 $\pm$ 13 19 $\pm$ 1.8 Large Top Left 72 $\pm$ 10 33 $\pm$ 2.9 52 $\pm$ 3.3 22 $\pm$ 4.2 300 $\pm$ 0 300 $\pm$ 0.0 21 $\pm$ 2.8 Top Right 220 $\pm$ 20 49 $\pm$ 7.7 220 $\pm$ 28 190 $\pm$ 27 280 $\pm$ 20 110 $\pm$ 36 27 $\pm$ 2.6 Bottom Right 280 $\pm$ 15 34 $\pm$ 1.8 160 $\pm$ 22 140 $\pm$ 22 280 $\pm$ 21 260 $\pm$ 19 21 $\pm$ 1.8 Top Center 220 $\pm$ 28 48 $\pm$ 5.2 120 $\pm$ 8.8 59 $\pm$ 12 240 $\pm$ 23 33 $\pm$ 8.5 39 $\pm$ 6.2 Aggregated 200 $\pm$ 8.9 41 $\pm$ 2.7 140 $\pm$ 13 100 $\pm$ 13 270 $\pm$ 12 180 $\pm$ 13 27 $\pm$ 1.7 Ultra Top Left 76 $\pm$ 7.0 34 $\pm$ 4.9 91 $\pm$ 11 36 $\pm$ 11 39 $\pm$ 21 15 $\pm$ 5.3 17 $\pm$ 3.6 Top Right 300 $\pm$ 0.0 92 $\pm$ 20 290 $\pm$ 7.8 120 $\pm$ 14 260 $\pm$ 19 150 $\pm$ 32 37 $\pm$ 5.5 Bottom Right 300 $\pm$ 0.0 70 $\pm$ 8.0 300 $\pm$ 0.0 130 $\pm$ 16 240 $\pm$ 28 67 $\pm$ 12 34 $\pm$ 6.0 Top Center 230 $\pm$ 35 29 $\pm$ 5.5 230 $\pm$ 29 75 $\pm$ 16 100 $\pm$ 32 17 $\pm$ 1.7 22 $\pm$ 4.4 Aggregated 230 $\pm$ 9.3 56 $\pm$ 5.4 230 $\pm$ 9.1 90 $\pm$ 7.1 160 $\pm$ 14 61 $\pm$ 8.2 27 $\pm$ 2.4 Aggregated 190 $\pm$ 6.9 46 $\pm$ 2.3 160 $\pm$ 6.3 80 $\pm$ 5.9 210 $\pm$ 11 130 $\pm$ 7.4 25 $\pm$ 1.4 Table 1: The number of environment steps (×$\times 10^{3}$) taken before the agent find the goal. Lower is better. The first goal time is considered to be $300\times 10^{3}$ steps if the agent never finds the goal. We see that our method is the most consistent, achieving performance as good as or better than all other methods in each of the 4 goals across 3 different maze layouts. The error quantity indicated is standard error over 8 seeds.
5.5 Does offline data help our method to compose skills better for faster exploration?
Figure 5 shows the percentage of the maze that the agent has covered throughout the training. The coverage of skill-based methods that do not use prior data during online learning, Online w/ Trajectory Skills and Online w/ HILP Skills, significantly lags behind baselines that use offline data after 50,000 environment steps. Many methods achieve similar coverage on antmaze-medium, likely because the maze is too small to differentiate the different methods. Ours is able to achieve the highest coverage on the antmaze-ultra, and is only surpassed on antmaze-large by HILP w/ Offline Data, which has high first goal times and slow learning. Thus, the coverage difference can likely be at least partially attributed to HILP w/ Offline Data struggling to find the goal and continuing to explore after finding the goal. All non-skill based methods struggle to get competitive coverage levels on antmaze-large and antmaze-ultra. This suggests both pretraining skills and the ability to leverage prior data online are crucial for efficient exploration, and our method effectively compounds their benefits. For completeness, we include the success rate and coverage results for each maze layout and goal location in Appendix F.
@@ Appendix A Compute Resources
We run all our experiments on NVIDIA A5000 GPUs.
@@
Thus, in total the results in this paper required approximately $13616$ GPU hours, or about 1.5 GPU years. Note this is an approximate upper bound, since not all methods required training checkpoints, and checkpoints were shared between different baselines that both uses trajectory skills or both used HILP skills.
@@ Appendix B VAE Architecture and hyperparameters
We use a VAE implementation from Park et al. (2024). The authors kindly shared with us their OPAL implementation (which produces the results of the OPAL baseline in the paper). In this implementation, the VAE encoder is a recurrent neural network that uses gated-recurrent units (Cho et al., 2014) (GRU). It takes in a short sequence of states and actions, and produces a probabilistic output of a latent $z$. The reconstruction policy decoder is a fully-connected network with ReLU activation (Nair & Hinton, 2010) that takes in both the state in the sequence as well as the latent $z$ to output an action distribution.
@@
GRU Hidden Size 256
GRU Layers 2 hidden layers
KL Coefficient ($\beta)$ 0.1
@@
In the online phase, our high-level policy is a Soft-Actor-Critic (SAC) agent (Haarnoja et al., 2018) with 10 critic networks, entropy backup disabled and LayerNorm added to the critics following the architecture design used in RLPD (Ball et al., 2023). We follow a similar strategy in ExPLORe (Li et al., 2024) where we sample 128 offline samples and 128 online samples and add RND reward bonus to all of the samples. The main difference is in the original ExPLORe paper is that they only add reward bonus to the offline data as additionally adding the bonus to the online replay buffer does not help for the maze goals they tested. In our experiments, we add the reward bonus to both offline data and online data, as it leads to better performance in goals where there is limited offline data coverage (see Appendix F).
@@
Discount factor ($\gamma$) 0.99
@@
Critic minimum ensemble size 1 for all methods on AntMaze, 2 for all on Kitchen,
@@
UTD Ratio 20 for AntMaze and Kitchen, 40 for Visual AntMaze
@@
Network Width 256
@@
Initial Entropy Temperature 1.0 on Kitchen, 0.05 on Visual AntMaze/AntMaze
@@ Appendix C Implementation details for baselines
Diffusion BC + JSRL. We use the diffusion model implementation from (Hansen-Estruch et al., 2023). Following the paper’s implementation, we train the model for 3 million gradient steps with a dropout rate of 0.1 and a cosine decaying learning rate schedule from the learning rate of $0.0003$. In the online phase, in the beginning of every episode, with probability $p$, we rollout the diffusion policy for a random number of steps that follows a geometric distribution $\mathrm{Geom}(1-\gamma)$ before sampling actions from the online agent (inspired by (Li et al., 2023)). A RND bonus is also added to the online batch on the fly with a coefficient of $2.0$ to further encourage online exploration of the SAC agent. The same coefficient is used in all other non-skill based baselines. For skill based baselines, we scale up the RND coefficient by the horizon length (4) to account for different reward scale. Following the BC + JSRL baseline used in ExPLORe (Li et al., 2024), we use a SAC agent with an ensemble of 10 critic networks, one actor network, with no entropy backup and LayerNorm in the critic networks. This configuration is used for all baselines on all environments. On AntMaze, We perform a hyperparameter sweep on both $p=\{0.5,0.75,0.9\}$ and the geometric distribution parameter $\gamma=\{0.99,0.995,0.997\}$ on the large maze with the top right goal and find that $p=0.9$ and $\gamma=0.99$ works the best. We also use these parameters for the Visual AntMaze experiments. For Kitchen, we perform a sweep on the parameter $p=\{0.2,0.5,0.75,0.9\}$ and find that 0.75 works best. We still use $\gamma=0.99$. We take the minimum of one random critic for AntMaze and Visual AntMaze, and the minimum of two random critics for Kitchen. For all methods, we use the same image encoder used in RLPD (Ball et al., 2023) for the Visual AntMaze task, with a latent dimension of 50 (encoded image is a 50 dimensional vector), which is then concatenated with proprioceptive state observations.
ExPLORe. We directly use the open-source implementation from https://github.com/facebookresearch/ExPLORe/. The only difference we make is to adjust the RND coefficient from $1.0$ to $2.0$ and additionally add such bonus to the online replay buffer (the original method only adds to the offline data). Empirically, we find a slightly higher RND coefficient improves performance slightly. Setting the value too high (e.g., 10) can in turn hurt performance. The SAC configuration is the same as that of the Diffusion BC + JSRL agent. We found that taking the minimum of one random critic on Visual AntMaze worked better for ExPLORe than taking the minimum of two, so all non-skill based baselines use this hyperparameter value.
@@
Online RL with HILP skills. This baseline is the same as the one above but with the skills from a recent unsupervised offline skill discovery method, HILP (Park et al., 2024). We use the official open-source implementation https://github.com/seohongpark/HILP and run the pretraining to obtain the skill policies. Then, we freeze the skill policies and learn a high-level RL agent to select skills every $H$ steps.
@@
where $\alpha$ controls the strength of the exploration tendency (RND coefficient). For AntMaze environments (both state-based and visual), we find that it is sufficient to use the minimum reward, $-1$, to label the offline data without a performance drop, so we opt for such a simpler design for our experiments.
@@
Appendix D ICVF Implementation Details and Ablation Experiments for Visual Antmaze

### v2

@@ 5.3 Can we leverage unsupervised trajectory skills to accelerate online learning?
Figure 3 shows the aggregated performance of our approach on all seven domains. Our method outperforms all prior methods on all domains except Scene, where our novel baseline HILP with Offline Data performs slightly better. It is worth noting that HILP with Offline Data also leverages offline data twice (one of the key ideas behind our method), both during offline and online learning. Both HILP-based methods (Online with HILP Skills and HILP with Offline Data) perform well on Scene, Single Cube, and AntSoccer, but struggle to learn on other tasks. The Online with Trajectory Skills baseline also consistently performs worse than our method across all seven domains, which demonstrates the importance of using prior data for online learning of the high-level policy, since that is only difference between this baseline and Ours. ExPLORe uses offline data during online learning, but does not pretrain skills, leading to slower learning on all seven environments and difficulty achieving any significant return on any domains other than the easier AntMaze and Single Cube tasks. We also report the performance on individual AntMaze mazes and Kitchen tasks in Figure 10, and observe that our method outperforms the baselines more on harder environments. This trend continues with HumanoidMaze, where individual results in Figure 15 show that Ours is the only method to achieve nonzero final return on the more difficult large and giant mazes. These experiments suggest that pretraining skills and the ability to leverage the prior data are both crucial for achieving efficient online learning, especially in more challenging environments. For the visual domain, we additionally perform an ablation study to assess the importance of the ICVF pre-trained representation, which we include in Appendix E. While ICVF combines synergistically with our method to further accelerate learning and exploration, initializing RND image encoder weights using ICVF is not critical to its success.
@@ 5.4 Is our method able to find goals faster than prior methods?
Even though we have demonstrated that our method is able to achieve a higher success rate faster than prior works, it is still not clear if our method actually leads to better exploration (instead of simply learning the high-level policy better). In this section, we study the exploration aspect in isolation in the AntMaze domain. Figure 4 reports the number of online environment interaction steps for the agent to reach the goal for the first time. Such a metric allows us to assess how efficiently the agent explores in the maze whereas the success rate metric only measures how good the agent is at reaching the desired goal. It is possible for an agent to be good at exploration, but bad at reaching goals consistently and vice-versa. Figure 4 shows that our method reaches goals faster than every baseline on all three maze layouts, which confirms that our method not only learns faster, but does so by exploring more efficiently. For completeness, Table 3 reports the same metric but for each individual goal location and maze layout, and we also include the success rate for each maze layout and goal location in Appendix H.2.
Figure 4: Interquartile mean (IQM) of the number of environment steps taken to reach the goal (smaller the better). The first goal time is considered to be $300\times 10^{3}$ steps if the agent never finds the goal. Our method is the most consistent, achieving performance better than all other baselines on all layouts (4 tasks/goals for each layout and 8 seeds for each task). The plot is generated using the rliable library with 95% stratified bootstrap confidence intervals (Agarwal et al., 2021).
5.5 How sensitive is the performance of our method is to its hyperparameters?
In this section, we study the sensitivity of two hyperparameters, 1) $\alpha$: the amount of optimism (RND coefficient) used when labeling offline data with UCB rewards (Equation 2), and 2) $H$: the length of the skill. We select one representative task from each domain and study the how different $\alpha$ and $H$ values affect our performance on these tasks (Figure 5). For the RND coefficient, we test a wide range of values. When the RND bonus is removed ($\alpha=0$), our method performs significantly worse, highlighting the importance of optimistic labeling on efficient online exploration. While we observe some variability across individual tasks (Appendix G, Figure 8), our method is generally not very sensitive to the RND coefficient value. The aggregated performance is similar for $\alpha\in\{2,8,16\}$. Another key hyperparmeter in our method is the skill horizon length (see Figure 5, right). We find that while there is some variability across individual tasks (Appendix G, Figure 9), a skill horizon length of 4 generally performs the best, and shorter or longer horizons perform much worse on certain tasks. We use $H=4$ for all experiments in the paper.
Figure 5: Sensitivity analysis on the RND coefficient ($\alpha$) and the skill horizon length ($H$) on a subset of tasks. We report the interquartile mean (IQM) of the normalized return across seven domains (we select one representative task per domain) for different hyperparameter values. The performance of our method is not very sensitive to the magnitude of $\alpha$ as long as it is within a reasonable range $(2,16)$. Without the bonus ($\alpha=0$), our method performs significantly worse. A skill horizon length of 4 performs significantly better than a horizon length of 2 or 8. We use a skill length of 4 and $\alpha=8$ for all experiments with skill-based methods. The normalized return for each individual task we use for this analysis can be found in Appendix G.
@@ Appendix A Compute Resources
We run all our experiments on NVIDIA A5000 GPU and V100 GPUs.
@@
[…]
Thus, in total the results in this paper required approximately $16,601$ GPU hours, or about 1.9 GPU years. Note this is an approximate upper bound, since not all methods required training checkpoints, and checkpoints were shared between different baselines that both uses trajectory skills or both used HILP skills.
@@ Appendix B VAE Architecture and hyperparameters
We use a VAE implementation from Park et al. (2024b). The authors kindly shared with us their OPAL implementation (which produces the results of the OPAL baseline in the paper). In this implementation, the VAE encoder is a recurrent neural network that uses gated-recurrent units (Cho et al., 2014) (GRU). It takes in a short sequence of states and actions, and produces a probabilistic output of a latent $z$. The reconstruction policy decoder is a fully-connected network with ReLU activation (Nair & Hinton, 2010) that takes in both the state in the sequence as well as the latent $z$ to output an action distribution.
@@
GRU Hidden Size 256 (AntMaze, Kitchen, VisualAntMaze)
512 (AntSoccer, HumanoidMaze, Scene, Cube)
GRU Layers 2 hidden layers (AntMaze, Kitchen, VisualAntMaze)
3 hidden layers (AntSoccer, HumanoidMaze, Scene, Cube)
KL Coefficient ($\beta)$ 0.1 (AntMaze, HumanoidMaze, Kitchen, VisualAntmaze, AntSoccer)
0.2 (Cube, Scene)
@@
In the online phase, our high-level policy is a Soft-Actor-Critic (SAC) agent (Haarnoja et al., 2018) with 10 critic networks, entropy backup disabled and LayerNorm added to the critics following the architecture design used in RLPD (Ball et al., 2023). We follow a similar strategy in ExPLORe (Li et al., 2024) where we sample 128 offline samples and 128 online samples and add RND reward bonus to all of the samples. The main difference is in the original ExPLORe paper is that they only add reward bonus to the offline data as additionally adding the bonus to the online replay buffer does not help for the maze goals they tested. In our experiments, we add the reward bonus to both offline data and online data, as it leads to better performance in goals where there is limited offline data coverage (see Appendix H.2).
@@
Discount factor ($\gamma$) 0.99 (AntMaze, VisualAntmaze, Kitchen)
0.995 (HumanoidMaze, Cube, Scene, AntSoccer)
@@
Critic minimum ensemble size 1 for all methods on AntMaze HumanoidMaze, AntSoccer;
2 for all methods on Kitchen, Scene, Cube;
@@
UTD Ratio 20 for state-based domains, 40 for Visual AntMaze
@@
Network Width 256 (AntMaze, Kitchen, Visual AntMaze)
512 (HumanoidMaze, AntSoccer, Cube, Scene)
@@
Initial Entropy Temperature 1.0 on Kitchen, 0.05 on all other environments
@@ Appendix C Implementation details for baselines
Diffusion BC + JSRL. We use the diffusion model implementation from (Hansen-Estruch et al., 2023). Following the paper’s implementation, we train the model for 3 million gradient steps with a dropout rate of 0.1 and a cosine decaying learning rate schedule from the learning rate of $0.0003$. In the online phase, in the beginning of every episode, with probability $p$, we rollout the diffusion policy for a random number of steps that follows a geometric distribution $\mathrm{Geom}(1-\gamma)$ before sampling actions from the online agent (inspired by (Li et al., 2023)). A RND bonus is also added to the online batch on the fly with a coefficient of $2.0$ to further encourage online exploration of the SAC agent. The same coefficient is used in all other non-skill based baselines. For skill based baselines, we scale up the RND coefficient by the horizon length (4) to account for different reward scale. Following the BC + JSRL baseline used in ExPLORe (Li et al., 2024), we use a SAC agent with an ensemble of 10 critic networks, one actor network, with no entropy backup and LayerNorm in the critic networks. This configuration is used for all baselines on all environments. On AntMaze, We perform a hyperparameter sweep on both $p=\{0.5,0.75,0.9\}$ and the geometric distribution parameter $\gamma=\{0.99,0.995,0.997\}$ on the large maze with the top right goal and find that $p=0.9$ and $\gamma=0.99$ works the best. We also use these parameters for the Visual AntMaze experiments. On AntSoccer, we use the same $p$ but raise the discount $\gamma$ to match the environment discount rate of $0.995$. On HumanoidMaze, we perform a sweep over $p=\{0.5,0.75,0.9\}$ on humanoidmaze-medium-navigate-v0 and find that $p=0.75$ works best. We still use $\gamma=0.995$. On Scene and Cube, we do a similar sweep for $p$ on the first task of Scene, Single Cube, and Double Cube, and find that $p=0.5$ works best. We still use $\gamma=0.995$. For Kitchen, we perform a sweep on the parameter $p=\{0.2,0.5,0.75,0.9\}$ and find that 0.75 works best. We use $\gamma=0.99$. We take the minimum of one random critic for AntMaze, Visual AntMaze, AntSoccer, and HumanoidMaze, and the minimum of two random critics for Kitchen, Scene, and Cube. For all methods, we use the same image encoder used in RLPD (Ball et al., 2023) for the Visual AntMaze task, with a latent dimension of 50 (encoded image is a 50 dimensional vector), which is then concatenated with proprioceptive state observations.
ExPLORe. We directly use the open-source implementation from https://github.com/facebookresearch/ExPLORe/. The only difference we make is to adjust the RND coefficient from $1.0$ to $2.0$ and additionally add such bonus to the online replay buffer (the original method only adds to the offline data). Empirically, we find a slightly higher RND coefficient improves performance slightly. The SAC configuration is the same as that of the Diffusion BC + JSRL agent. We found that taking the minimum of one random critic on Visual AntMaze worked better for ExPLORe than taking the minimum of two, so all non-skill based baselines use this hyperparameter value.
@@
Online RL with HILP skills. This baseline is the same as the one above but with the skills from a recent unsupervised offline skill discovery method, HILP (Park et al., 2024b). We use the official open-source implementation https://github.com/seohongpark/HILP and run the pretraining to obtain the skill policies. Then, we freeze the skill policies and learn a high-level RL agent to select skills every $H$ steps.
@@
where $\alpha$ controls the strength of the exploration tendency (RND coefficient). For every environment with $-1$/00 rewards, we find that it is sufficient to use the minimum reward, $-1$, to label the offline data without a performance drop, so we opt for such a simpler design for our experiments. On Kitchen, we use a reward prediction model since the reward is the number of tasks currently completed.
@@
Appendix D Domain Details
We also include four additional domains repurposed from a goal-conditioned offline RL benchmark (Park et al., 2024a).
OGBench HumanoidMaze is a navigation task similar to AntMaze, but with the Ant agent replaced by a Humanoid agent. HumanoidMaze is much more challenging than AntMaze because Humanoid control is much harder with much higher action dimensionality (21 vs. 8). We benchmark on all six datasets in the benchmark. They involve three mazes of increasing size, humanoid-medium, humanoid-large, and humanoid-giant, and two types of datasets, navigate (collected by noisy expert policy that randomly navigates around the maze) and stitch (containing only short segments that test the algorithm’s ability to stitch them together).
OGBench AntSoccer is a navigation task similar to AntMaze, but with the added complexity where the Ant agent must first travel to the location of a soccer ball, then dribble the soccer ball to the goal location. We benchmark on the two navigate datasets, antsoccer-arena-navigate and antsoccer-medium-navigate.
OGBench Cube and Scene are two manipulation domains with a range of manipulation tasks for each domain. Cube involves using a robot arm to arrange blocks from an initial state to a goal state, which requires pick and place actions to move and/or stack blocks. We benchmark on the two sub-domains, cube-single and cube-double which include one and two blocks, respectively. We also benchmark on the Scene environment with two buttons, a drawer, and a window. The most difficult task requires the agent to perform 8 atomic actions: unlock the drawer and window, open drawer, pick up block, place block in drawer, close drawer, open window, lock drawer and window. We benchmark on all 5 tasks for cube-single, cube-double, and scene, and use the play datasets collected by non-Markovian expert policies with temporally correlated noise.
Aside from the state-based domains above, we also consider a visual domain below to test the ability of our method in scaling up to high-dimensional image observations.
Appendix E ICVF Implementation Details and Ablation Experiments for Visual Antmaze